I need help with Quadratic Equations. I don't know how to solve these problem.
2Z^2+3=341
(p+6)^2=3
p^2+5p-5=0

Answer by njusa(12)

2Z^2+3=341
2Z^2= 348
divide by 2
Z^2 = 174
z= + or - sqrt(174)
(P+6)^2 = 3
(p+6)(p+6) = 3
p^2 +12p+36 = 3
p^2 +12p+33=0
(-12 +- sqrt(144 -4*33))/2
-12+-2root(3)/2
-6 +- root(3)
use the same method for the last one